OpenStudy (anonymous):

4x^2+8x-11x+6-5x^2+9 A -x^2-3x+8 B -x^2+3x+3 C x^2-3x+8 D x^2+3x+3

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

i think its A or B

OpenStudy (anonymous):

wrong equation

OpenStudy (kitten_is_back):

sorry try tiger aglba

OpenStudy (anonymous):

4x^2+8x-11x+6-5x^2+2 A -x^2-3x+8------ B -x^2+3x+3 C x^2-3x+8------ D x^2+3x+3

OpenStudy (anonymous):

omg B and D those 3 at the end is suppose to be eight

OpenStudy (kitten_is_back):

http://www.tiger-algebra.com/drill/4x~2_8x-11x_6-5x~2_2/

OpenStudy (anonymous):

4x^2+8x-11x+6-5x^2+2 A -x^2-3x+8------ B -x^2+3x+8 C x^2-3x+8------ D x^2+3x+8

OpenStudy (kitten_is_back):

logo tiger Home | Latest | Discuss | About | Help | Translation 8,696,210 solved | 670 online 4x2+8x-11x+6-5x2+2 Final result : -1 • (x2 + 3x - 8) Step by step solution : Step 1 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 1 : (((((4•(x2))+8x)-11x)+6)-(5•x2))+2 Step 2 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 2 : (((((4•x2)+8x)-11x)+6)-5x2)+2 Step 3 : Simplify -x2-3x+6 + 2 Pulling out like terms : 3.1 Pull out like factors : -x2 - 3x + 8 = -1 • (x2 + 3x - 8) Trying to factor by splitting the middle term 3.2 Factoring x2 + 3x - 8 The first term is, x2 its coefficient is 1 . The middle term is, +3x its coefficient is 3 . The last term, "the constant", is -8 Step-1 : Multiply the coefficient of the first term by the constant 1 • -8 = -8 Step-2 : Find two factors of -8 whose sum equals the coefficient of the middle term, which is 3 . -8 + 1 = -7 -4 + 2 = -2 -2 + 4 = 2 -1 + 8 = 7 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Final result : -1 • (x2 + 3x - 8)

OpenStudy (anonymous):

i think its A